# Leonardo V. Santoro   Victor M. Panaretos

### v1

@@ Abstract
We propose a novel kernel-based nonparametric two-sample test, employing the combined use of kernel mean and kernel covariance embedding. Our test builds on recent results showing how such combined embeddings map distinct probability measures to mutually singular Gaussian measures on the kernel’s RKHS. Leveraging this result, we construct a test statistic based on the relative entropy between the Gaussian embeddings, i.e. the likelihood ratio. The likelihood ratio is specifically tailored to detect equality versus singularity of two Gaussians, and satisfies a $0/\infty$” law, in that it vanishes under the null and diverges under the alternative. To implement the test in finite samples, we introduce a regularised version, calibrated by way of permutation. We prove consistency, establish uniform power guarantees under mild conditions, and discuss how our framework unifies and extends prior approaches based on spectrally regularized MMD. Empirical results on synthetic and real data demonstrate remarkable gains in power compared to state-of-the-art methods, particularly in high-dimensional and weak-signal regimes.
@@ 1 Introduction
based on i.i.d. random samples $\{X_{1},\dots,X_{n}\}\sim\mathbb{P}$ and $\{Y_{1},\dots,Y_{m}\}\sim\mathbb{Q}$, for $n,m\in\mathbb{N}$, without making parametric assumptions on either $\mathbb{P}$ or $\mathbb{Q}$. Nonparametric two-sample testing is a fundamental problem in statistical inference, dating back to the foundational work of Pearson [33] on the $\chi^{2}$-tests and $t$-tests. Early work on one-dimensional distributions made use of the order structure of the real line, using the empirical distribution function (as in the Kolmogorov–Smirnov test [45]), or rank information (Wilcoxon [52] and Mann–Whitney [28, 26]). In higher dimensions, the problem is considerably more challenging: the vague specification of the alternative allows for manifold and subtle differences –possibly exacerbated by complexity/dimensionality– and does not inform on how to design efficient criteria. And yet, nonparametric tests are especially well-suited to high-dimensional and complex settings, where the researcher is either unable to or unwilling to commit to a strong specification of the hypothesis pair.
@@
Recent work by Santoro et al. [40] has provided a new theoretical perspective. They show that when probability measures are embedded into an RKHS via their combined kernel mean and covariance embeddings, they can be identified with Gaussian measures $\mathcal{N}_{\mathbb{P}}$ and $\mathcal{N}_{\mathbb{Q}}$ that either coincide or are mutually singular. That is, the combined embedding translates $\{H_{0}:\mathbb{P}=\mathbb{Q}\mbox{ vs }H_{1}:\mathbb{P}\neq\mathbb{Q}\}$ to the considerably sharper pair $\{H_{0}:\mathcal{N}_{\mathbb{P}}=\mathcal{N}_{\mathbb{Q}}\mbox{ vs }H_{1}:\mathcal{N}_{\mathbb{P}}\perp\mathcal{N}_{\mathbb{Q}}\}$. From an information-theoretic perspective, singular measures are infinitely different, showing that even subtle differences between $\mathbb{P}$ and $\mathbb{Q}$ are greatly amplified in the RKHS: while the original hypothesis test may involve subtle differences between distributions, in the kernel-embedded space, the alternative hypothesis is always maximally separated from the null. This embedding perspective not only sheds light on the reasons for the remarkable empirical performance of kernel methods, but also suggests that the performance can be enhanced considerably further, as existing strategies do not explicitly target the singularity phenomenon (indeed, Santoro et al. [40] show that existing methods may well fail to harness it depending on the scenario). The purpose of this paper is to precisely make full use of the potential offered by the singularity phenomenon, by specifically employing information-theoretic quantities targeting it. In particular, the Kullback–Leibler (KL) divergence between the embedded Gaussian measures is either zero (under $H_{0}$) or infinite (under $H_{1}$). This provides a natural basis for constructing a powerful test statistic tailored to the very sharp theoretical separation between the null and alternative.
@@
We propose a new kernel-based two-sample test built on the embedded Gaussians’ likelihood ratio (equivalently, relative entropy). This idea is implemented by comparing the kernel embeddings of the empirical distributions $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ through the suitably regularised relative entropy of their induced empirical Gaussian embeddings $\mathcal{N}_{\mathbb{P}_{n}}$ and $\mathcal{N}_{\mathbb{Q}_{n}}$. Regularisation is called for to induce stability at the finite sample level. The test statistic is simple to compute, yet grounded in concrete geometric and probabilistic insights. Unlike previous approaches where the criterion depends smoothly on the magnitude of discrepancies between $\mathbb{P}$ and $\mathbb{Q}$, our criterion enjoys a much sharper dichotomy: at the population level, any departure from the null leads to asymptotically maximal separation. Concretely, we show that the regularised test statistic is finite for all values of the regularisation parameter, vanishes under the null, and diverges under the alternative as the regularisation parameter vanishes (Theorem 3.3), a dichotomy crystallising in the 0–1 law given by Corollary 3.9. We establish theoretical guarantees for the proposed test, including consistency against fixed alternatives (Theorem 3.10) and uniform consistency under a suitable notion of separation between $\mathbb{P}$ and $\mathbb{Q}$ (Theorem 3.11), for translation-invariant kernels. These results hold under mild conditions and provide a precise characterization of the regularization rate required for consistency. In addition, our framework offers a unifying perspective on previously proposed kernel-based tests. In particular, it shows that spectral regularization methods [8, 16] are a restricted version of a broader information-theoretic criterion, shedding light on the mechanisms that govern their effectiveness. We illustrate the finite-sample performance of our test on a range of synthetic and real datasets. We empirically find that the method offers substantial power improvements compared to existing approaches, excelling in challenging scenarios involving subtle and/or high-dimensional differences between distributions.
@@ 2.1 Reproducing Kernel Hilbert Spaces
The KFD thus seeks the direction that maximizes the mean difference $\mu_{\mathbb{P}}-\mu_{\mathbb{Q}}$ taking into account the average variability under $\mathbb{P}$ and $\mathbb{Q}$, and can be seen as a refinement of MMD that adapts the testing direction according to the geometry of the two distributions. In a similar spirit, [16] generalise this approach and proposes a family of two-sample tests based on spectrally regularized differences between kernel mean embeddings, including but not limited to
@@
which incorporates the spectral decay of the covariance into the mean difference in a variety of possible ways. [16] further show that with appropriate choice of the regularization, the spectral regularised statistic is consistent and can achieve minimax optimality over suuitably separated alternatives.
@@ 3 Regularised Kernel Likelihood Testing
Let $\mathbb{P},\mathbb{Q}$ be probability measures on a compact and separable metric space $\mathcal{X}$, and $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ be a universal reproducing kernel. Then:
@@
$\mathcal{N}(\mathbf{m}_{1},\mathbf{C}_{1})\sim\mathcal{N}(m_{2},\mathbf{C}_{2})\quad\Longleftrightarrow\quad\begin{cases}\mathbf{m}_{2}-\mathbf{m}_{1}\in\operatorname{Range}(\bar{\mathbf{C}}^{\nicefrac{{1}}{{2}}}),\quad\text{where }\bar{\mathbf{C}}=\tfrac{1}{2}(\mathbf{C}_{1}+\mathbf{C}_{2}),\quad\text{and}\\ \mathbf{C}_{1}=\mathbf{C}_{2}^{\nicefrac{{1}}{{2}}}(\mathbf{H}-\mathbf{I})\mathbf{C}_{2}^{\nicefrac{{1}}{{2}}},\quad\text{for some Hilbert-Schmidt }\mathbf{H}\succ-\mathbf{I}.\end{cases}$
@@
Consider probability distributions $\mu,\nu$ on $\mathcal{H}$, and suppose  $\mu\ll\nu$, so that the Radon-Nikodyn derivative $d\mu/d\nu$ exists $\nu$-almost everywhere, in that $\mu(\mathcal{A})=\int_{\mathcal{A}}\left({d\mu}/{d\nu}\right)(x)d\nu(x)$ for any measurable set $\mathcal{A}\subset\mathcal{X}$. With this notation, when $\mu\ll\nu$, the Kullback-Leibler divergence of $\mu$ wrt to $\nu$, or relative entropy, is given by $D_{\mathrm{KL}}(\mu,\nu)=\int_{\mathcal{X}}\log\left({d\mu}/{d\nu}\right)d\mu.$ Focusing on the case of (equivalent) Gaussian measures $\mu=\mathcal{N}(\mathbf{m}_{1},\mathbf{S}_{1})$ and $\nu=\mathcal{N}(\mathbf{m}_{2},\mathbf{S}_{2})$, the Kullback-Leibler divergence admits a particularly tractable form:
@@
$\begin{split}D_{\gamma,\mathrm{KL}}\big{(}\mathcal{N}(\mathbf{m}_{\mathbb{Q}},\mathbf{S}_{\mathbb{P}})\>||\>\mathcal{N}(\mathbf{m}_{\mathbb{P}},\mathbf{S}_{\mathbb{Q}})\big{)}:=&\frac{1}{2}\|(\mathbf{S}_{\mathbb{Q}}+\gamma\mathbf{I})^{-\nicefrac{{1}}{{2}}}(\mathbf{m}_{\mathbb{P}}-\mathbf{m}_{\mathbb{Q}})\|_{\mathcal{H}}^{2}\\ &\>-\frac{1}{2}\log\operatorname{det}_{2}\left(\mathbf{I}+(\mathbf{S}_{\mathbb{Q}}+\gamma\mathbf{I})^{-\nicefrac{{1}}{{2}}}(\mathbf{S}_{\mathbb{P}}-\mathbf{S}_{\mathbb{Q}})(\mathbf{S}_{\mathbb{Q}}+\gamma\mathbf{I})^{-\nicefrac{{1}}{{2}}}\right).\end{split}$ (5)
@@
$\mathfrak{T}_{\gamma}(\mathbb{P},\mathbb{Q}):=D_{\gamma,\mathrm{KL}}\big{(}\mathcal{N}(\mathbf{m}_{\mathbb{Q}},\mathbf{S}_{\mathbb{P}})\>||\>\mathcal{N}(\mathbf{m}_{\mathbb{P}},\mathbf{S}_{\mathbb{Q}})\big{)}$ (6)
@@
Let $\mathbb{P},\mathbb{Q}$ be probability measures on $\mathcal{X}$. Let $d\geq 1$ and $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ be a universal reproducing kernel. Then $\mathfrak{T}_{\gamma}(\mathbb{P},\mathbb{Q})$ is finite for any positive $\gamma>0$, while:
@@ 3.1 The Two-Sample Test
The critical value $\tau_{nm}$ will be chosen by way of a permutation method. For given values $n,m\geq 0$, let $\{Z_{j}\}_{j=1}^{n+m}$ denote the pooled sample $\{X_{1},\dots,X_{n},Y_{1},\dots,Y_{m}\}$. Let $\sigma$ be a uniformly random permutation of $\{1,\dots,n+m\}$, denoted $\sigma\in\mathfrak{S}_{n+m}$. The $(1-\alpha)$-quantile of the permutation distribution of the regularized test statistic as:
@@ 3.3 Implementation
As discussed in Section 3.1, the test statistic will be calibrated by way of permutation: we considers the distribution of the test statistic computed over all label permutations of the pooled sample,
@@
where $\mathfrak{S}_{n+m}$ is the symmetric group of permutations of $n+m]=\{1,\dots,n+m\}$. In practice, we approximate this distribution is approximated by sampling $B$ permutations $\sigma_{1},\dots,\sigma_{B}$ uniformly at random from $\mathfrak{S}_{n+m}$, and computing the empirical $(1-\alpha)$-quantile of the resulting values. This sample yields an approximation $\hat{q}_{1-\alpha}^{B,\gamma}$ to the quantile defined in (11), namely
@@
    The choice of regularization strength $\gamma$ regulates a tradeoff between finite-sample stability and making use of the singularity phenomenon. To ensure maximal gains in power, our proposal is to consider testing across a range of values $\gamma$, and adjusting for multiple testing. This is reminiscent of the idea of adapting to a family of kernels, that has been explored in regression and classification settings under the name multiple-kernel learning [13]. This approach is seen to yield very good performance – the gains in power to be made by adaptive choice seem to clearly outweigh the cost of the multiple testing correction.
@@ 4.2 Real Data
The MNIST dataset contains 70,000 handwritten digit images [24]. We compare the distribution $\mathbb{P}$ of samples of true MNIST images to a distribution $\mathbb{Q}$ of samples by a pretrained DCGAN [35]. Samples from both distributions are shown in Figure 13. This setting mimics a transfer learning scenario where a model trained on real data is exposed to synthetic inputs at deployment, and is thus relevant to generative model evaluation under distribution shift.
For each sample size $n\in\{100,200,300\}$, we draw $N_{\text{iters}}=75$ independent pairs of samples $(X_{1},\dots,X_{n})\sim\mathbb{P}$ and $(Y_{1},\dots,Y_{n})\sim\mathbb{Q}$. Each run produces a $p$-value computed from $200$ permutations, and the reported rejection rate is the proportion of the 75 repetitions in which the null hypothesis was rejected. To assess type I error control, we repeat the same procedure with both samples drawn from the pooled data, verifying that the empirical rejection rates are close to the nominal level. Samples from both distributions are shown in Figure 13, along with the average rejection rates for the three sample sizes, under the alternative (drawing samples from MNIST and fake MNIST datasets, respectively) and null (drawing both samples from the mixture).
@@
For each sample size $n\in\{250,500,750,1000\}$, we draw $N_{\text{iters}}=75$ independent pairs of samples $(X_{1},\dots,X_{n})\sim\mathbb{P}$ and $(Y_{1},\dots,Y_{n})\sim\mathbb{Q}$. Each run produces a p-value computed from $B=200$ permutations, and the reported rejection rate is the proportion of the 75 repetitions in which the null hypothesis was rejected. To assess type I error control, we repeat the same procedure with both samples drawn from the pooled data, verifying that the empirical rejection rates are close to the nominal level. Samples from both distributions are shown in Figure 14, along with the average rejection rates for the three sample sizes, under the alternative (drawing samples from CIFAR-10 and CIFAR-10.1 datasets, respectively) and null (drawing both samples from the mixture).

### v2

@@ Abstract
We propose a novel kernel-based nonparametric two-sample test, employing the combined use of kernel mean and kernel covariance embedding. Our test builds on recent results showing how such combined embeddings map distinct probability measures to mutually singular Gaussian measures on the kernel’s RKHS. Leveraging this “separation of measure phenomenon”, we construct a test statistic based on the relative entropy between the Gaussian embeddings, in effect the likelihood ratio. The likelihood ratio is specifically tailored to detect equality versus singularity of two Gaussians, and satisfies a $0/\infty$” law, in that it vanishes under the null and diverges under the alternative. To implement the test in finite samples, we introduce a regularised version, calibrated by way of permutation. We prove consistency, establish uniform power guarantees under mild conditions, and discuss how our framework unifies and extends prior approaches based on spectrally regularized MMD. Empirical results on synthetic and real data demonstrate remarkable gains in power compared to state-of-the-art methods, particularly in high-dimensional and weak-signal regimes.
@@ 1 Introduction
based on i.i.d. random samples $\{X_{1},\dots,X_{n}\}\sim\mathbb{P}$ and $\{Y_{1},\dots,Y_{m}\}\sim\mathbb{Q}$, for $n,m\in\mathbb{N}$, without making parametric assumptions on either $\mathbb{P}$ or $\mathbb{Q}$. Nonparametric two-sample testing is a fundamental problem in statistical inference, dating back to the foundational work of Pearson [33] on the $\chi^{2}$-test. Early work on one-dimensional distributions made use of the order structure of the real line, using the empirical distribution function (as in the Kolmogorov–Smirnov test [45]), or rank information (Wilcoxon [52] and Mann–Whitney [28, 26]). In higher dimensions, the problem is considerably more challenging: the vague specification of the alternative allows for manifold and subtle differences –possibly exacerbated by complexity/dimensionality– and does not inform on how to design efficient criteria. And yet, nonparametric tests are especially well-suited to high-dimensional and complex settings, where the researcher is either unable or unwilling to commit to a strong specification of the hypothesis pair.
@@
Recent work by Santoro et al. [40] has provided a new theoretical perspective by eliciting a “separation of measure phenomenon”. They show that when distinct probability measures are embedded into an RKHS via their combined kernel mean and covariance embeddings, they can be identified with Gaussian measures $\mathcal{N}_{\mathbb{P}}$ and $\mathcal{N}_{\mathbb{Q}}$ that are mutually singular. That is, the combined embedding translates $\{H_{0}:\mathbb{P}=\mathbb{Q}\mbox{ vs }H_{1}:\mathbb{P}\neq\mathbb{Q}\}$ to the considerably sharper pair $\{H_{0}:\mathcal{N}_{\mathbb{P}}=\mathcal{N}_{\mathbb{Q}}\mbox{ vs }H_{1}:\mathcal{N}_{\mathbb{P}}\perp\mathcal{N}_{\mathbb{Q}}\}$. From an information-theoretic perspective, singular measures are infinitely different, showing that even subtle differences between $\mathbb{P}$ and $\mathbb{Q}$ are greatly amplified in the RKHS: while the original hypothesis pair may be arbitrarily close, in the kernel-embedded space, the the hypotheses are maximally separated. This embedding perspective not only sheds light on the reasons for the remarkable empirical performance of kernel methods, but also suggests that the performance can be enhanced considerably further, as existing strategies do not explicitly target the separation of measure phenomenon (indeed, Santoro et al. [40] show that existing methods may well fail to harness it depending on the scenario). The purpose of this paper is to precisely make full use of the potential offered by this phenomenon, by specifically employing information-theoretic quantities targeting it. In particular, the Kullback–Leibler (KL) divergence between the embedded Gaussian measures is either zero (under $H_{0}$) or infinite (under $H_{1}$). This provides a natural basis for constructing a powerful test statistic tailored to the very sharp theoretical separation between the null and alternative.
@@
We propose a new kernel-based two-sample test built on the embedded Gaussians’ likelihood ratio (relative entropy, to be precise). This idea is implemented by comparing the kernel embeddings of the empirical distributions $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ through the suitably regularised relative entropy of their induced empirical Gaussian embeddings $\mathcal{N}_{\mathbb{P}_{n}}$ and $\mathcal{N}_{\mathbb{Q}_{n}}$. Regularisation is called for to induce stability at the finite sample level. The test statistic is simple to compute, yet grounded in concrete geometric and probabilistic insights. Unlike previous approaches where the criterion depends smoothly on the magnitude of discrepancies between $\mathbb{P}$ and $\mathbb{Q}$, our criterion enjoys a much sharper dichotomy: at the population level, any departure from the null leads to asymptotically maximal separation. Concretely, we show that the regularised test statistic is finite for all values of the regularisation parameter, vanishes under the null, and diverges under the alternative as the regularisation parameter vanishes (Theorem 3.3), a dichotomy crystallising in the 0–1 law given by Corollary 3.9. We establish theoretical guarantees for the proposed test, including consistency against fixed alternatives (Theorem 3.10) and uniform consistency under a suitable notion of separation between $\mathbb{P}$ and $\mathbb{Q}$ (Theorem 3.11), for translation-invariant kernels. These results hold under mild conditions and provide a precise characterization of the regularization rate required for consistency. In addition, our framework offers a unifying perspective on previously proposed kernel-based tests. In particular, it shows that spectral regularization methods [8, 16] are a restricted version of a broader information-theoretic criterion, shedding light on the mechanisms that govern their effectiveness. We illustrate the finite-sample performance of our test on a range of synthetic and real datasets. We empirically find that the method offers substantial power improvements compared to existing approaches, excelling in challenging scenarios involving subtle and/or high-dimensional differences between distributions.
@@ 2.1 Reproducing Kernel Hilbert Spaces
The KFD thus seeks the direction that maximizes the mean difference $\mu_{\mathbb{P}}-\mu_{\mathbb{Q}}$ taking into account the average variability under $\mathbb{P}$ and $\mathbb{Q}$, and can be seen as a refinement of MMD that adapts the testing direction according to the geometry of the two distributions. In a similar spirit, [16] generalise this approach and propose a family of two-sample tests based on spectrally regularized differences between kernel mean embeddings, including but not limited to
@@
which incorporates the spectral decay of the covariance into the mean difference in a variety of possible ways. [16] further show that with appropriate choice of the regularization, the spectral regularised statistic is consistent and can achieve minimax optimality over suitably separated alternatives.
@@ 3 Regularised Kernel Likelihood Testing
Let $\mathbb{P},\mathbb{Q}$ be probability measures on a compact and separable metric space $\mathcal{X}$, and $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ be a universal reproducing kernel. Then:
@@
$\mathcal{N}(\mathbf{m}_{1},\mathbf{C}_{1})\sim\mathcal{N}(\mathbf{m}_{2},\mathbf{C}_{2})\quad\Longleftrightarrow\quad\begin{cases}\mathbf{m}_{2}-\mathbf{m}_{1}\in\operatorname{Range}(\bar{\mathbf{C}}^{\nicefrac{{1}}{{2}}}),\quad\text{where }\bar{\mathbf{C}}=\tfrac{1}{2}(\mathbf{C}_{1}+\mathbf{C}_{2}),\quad\text{and}\\ \mathbf{C}_{1}=\mathbf{C}_{2}^{\nicefrac{{1}}{{2}}}(\mathbf{H}-\mathbf{I})\mathbf{C}_{2}^{\nicefrac{{1}}{{2}}},\quad\text{for some Hilbert-Schmidt }\mathbf{H}\succ-\mathbf{I}.\end{cases}$
@@
Consider probability distributions $\mu,\nu$ on $\mathcal{H}$, and suppose  $\mu\ll\nu$, so that the Radon-Nikodym derivative $d\mu/d\nu$ exists $\nu$-almost everywhere, in that $\mu(\mathcal{A})=\int_{\mathcal{A}}\left({d\mu}/{d\nu}\right)(x)d\nu(x)$ for any measurable set $\mathcal{A}\subset\mathcal{X}$. With this notation, when $\mu\ll\nu$, the Kullback-Leibler divergence of $\mu$ wrt to $\nu$, or relative entropy, is given by $D_{\mathrm{KL}}(\mu,\nu)=\int_{\mathcal{X}}\log\left({d\mu}/{d\nu}\right)d\mu.$ Focusing on the case of (equivalent) Gaussian measures $\mu=\mathcal{N}(\mathbf{m}_{1},\mathbf{S}_{1})$ and $\nu=\mathcal{N}(\mathbf{m}_{2},\mathbf{S}_{2})$, the Kullback-Leibler divergence admits a particularly tractable form:
@@
$\begin{split}D_{\gamma,\mathrm{KL}}\big{(}\mathcal{N}(\mathbf{m}_{\mathbb{P}},\mathbf{S}_{\mathbb{P}})\>||\>\mathcal{N}(\mathbf{m}_{\mathbb{Q}},\mathbf{S}_{\mathbb{Q}})\big{)}:=&\frac{1}{2}\|(\mathbf{S}_{\mathbb{Q}}+\gamma\mathbf{I})^{-\nicefrac{{1}}{{2}}}(\mathbf{m}_{\mathbb{P}}-\mathbf{m}_{\mathbb{Q}})\|_{\mathcal{H}}^{2}\\ &\>-\frac{1}{2}\log\operatorname{det}_{2}\left(\mathbf{I}+(\mathbf{S}_{\mathbb{Q}}+\gamma\mathbf{I})^{-\nicefrac{{1}}{{2}}}(\mathbf{S}_{\mathbb{P}}-\mathbf{S}_{\mathbb{Q}})(\mathbf{S}_{\mathbb{Q}}+\gamma\mathbf{I})^{-\nicefrac{{1}}{{2}}}\right).\end{split}$ (5)
@@
$\mathfrak{T}_{\gamma}(\mathbb{P},\mathbb{Q}):=D_{\gamma,\mathrm{KL}}\big{(}\mathcal{N}(\mathbf{m}_{\mathbb{P}},\mathbf{S}_{\mathbb{P}})\>||\>\mathcal{N}(\mathbf{m}_{\mathbb{Q}},\mathbf{S}_{\mathbb{Q}})\big{)}$ (6)
@@
Let $\mathbb{P},\mathbb{Q}$ be probability measures on $\mathcal{X}$. Let $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ be a universal reproducing kernel. Then $\mathfrak{T}_{\gamma}(\mathbb{P},\mathbb{Q})$ is finite for any positive $\gamma>0$, while:
@@ 3.1 The Two-Sample Test
The critical value $\tau_{nm}$ will be chosen by way of a permutation method. For given values $n,m\geq 0$, let $\{Z_{j}\}_{j=1}^{n+m}$ denote the pooled sample $\{X_{1},\dots,X_{n},Y_{1},\dots,Y_{m}\}$. Let $\sigma$ be a uniformly random permutation of $\{1,\dots,n+m\}$, denoted $\sigma\in\mathfrak{S}_{n+m}$. The $(1-\alpha)$-quantile of the permutation distribution of the regularized test statistic is defined as:
@@ 3.3 Implementation
As discussed in Section 3.1, the test statistic will be calibrated by way of permutation: we consider the distribution of the test statistic computed over all label permutations of the pooled sample,
@@
where $\mathfrak{S}_{n+m}$ is the symmetric group of permutations of $n+m]=\{1,\dots,n+m\}$. In practice, this distribution is approximated by sampling $B$ permutations $\sigma_{1},\dots,\sigma_{B}$ uniformly at random from $\mathfrak{S}_{n+m}$, and computing the empirical $(1-\alpha)$-quantile of the resulting values. This sample yields an approximation $\hat{q}_{1-\alpha}^{B,\gamma}$ to the quantile defined in (11), namely
@@
    The choice of regularization strength $\gamma$ regulates a tradeoff between finite-sample stability and making use of the separation of measure phenomenon. To ensure maximal gains in power, our proposal is to consider testing across a range of values $\gamma$, and adjusting for multiple testing. This is reminiscent of the idea of adapting to a family of kernels, that has been explored in regression and classification settings under the name multiple-kernel learning [13]. This approach is seen to yield very good performance – the gains in power to be made by adaptive choice seem to clearly outweigh the cost of the multiple testing correction.
@@ 4.2 Real Data
The MNIST dataset contains 70,000 handwritten digit images [24]. We compare the distribution $\mathbb{P}$ of samples of true MNIST images to a distribution $\mathbb{Q}$ of samples by a pretrained DCGAN [35], trained to mimic the generation of handwritten digit samples in MNIST. Samples from both distributions are shown in Figure 13. This setting mimics a transfer learning scenario where a model trained on real data is exposed to synthetic inputs at deployment, and is thus relevant to generative model evaluation under distribution shift.
For each sample size $n\in\{100,200,300\}$, we draw $N_{\text{iters}}=75$ independent pairs of samples $(X_{1},\dots,X_{n})\sim\mathbb{P}$ and $(Y_{1},\dots,Y_{n})\sim\mathbb{Q}$. Each run produces a $p$-value computed from $200$ permutations, and the reported rejection rate is the proportion of the 75 repetitions in which the null hypothesis was rejected. To assess type I error control, we repeat the same procedure with both samples drawn from the pooled data, verifying that the empirical rejection rates are close to the nominal level. Samples from both distributions are shown in Figure 13, along with the average rejection rates for the three sample sizes, under the alternative (drawing samples from MNIST and fake MNIST datasets, respectively) and null distribution (drawing both samples from MNIST) respectively.
@@
For each sample size $n\in\{250,500,750,1000\}$, we draw $N_{\text{iters}}=75$ independent pairs of samples $(X_{1},\dots,X_{n})\sim\mathbb{P}$ and $(Y_{1},\dots,Y_{n})\sim\mathbb{Q}$. Each run produces a p-value computed from $B=200$ permutations, and the reported rejection rate is the proportion of the 75 repetitions in which the null hypothesis was rejected. To assess type I error control, we repeat the same procedure with both samples drawn from the pooled data, verifying that the empirical rejection rates are close to the nominal level. Samples from both distributions are shown in Figure 14, along with the average rejection rates for the three sample sizes, under the alternative (drawing samples from CIFAR-10 and CIFAR-10.1 datasets, respectively) and null (drawing both samples from the CIFAR-10).